OpenStudy (anonymous):

is X always ^2 or does it vary for functions will give medal

OpenStudy (anonymous):

x in function don't have to be squared ex: f(x) = x+1

OpenStudy (anonymous):

y = 5x + 25 y = 5(8) + 25 y = 40 + 25 y = 65 where did the 40 come from